OpenStudy (kg1975):

Plz Help!!! @aloneS @jhonyy9 @Awolflover1 @quickstudent @brainzonly

alones (alones):

your question?

OpenStudy (kg1975):

\[(1+\frac{ 1 }{ 3 })^{2}-\frac{ 2 }{ 9 }\]

OpenStudy (mrhoola):

P. E. M. D. A. S. Start from there .

OpenStudy (kittiwitti1):

First, convert the 1 to thirds fraction format.

OpenStudy (kittiwitti1):

Okay. When do you get back?

OpenStudy (kittiwitti1):

. Jeez.

jhonyy9 (jhonyy9):

1 + 1/3 = ?

jhonyy9 (jhonyy9):

ok i know but the rule of OS dont give possibile giving directly right answers -

OpenStudy (kittiwitti1):

That's why I said "when do you get back" ... you read that right?

jhonyy9 (jhonyy9):

just teach the right way how can be solved

jhonyy9 (jhonyy9):

ok. i got it

OpenStudy (kittiwitti1):

Yeah. I don't give direct answers, I tell people to stop giving them.

jhonyy9 (jhonyy9):

ok. nice work,job this is the right way teaching math sure bye bye